how to find solution for:
1) 3x+2y-7=0
2) 1234x+1345y-786=0
i dont know the perfect solution for solving both the equation except hit and trial. but it won't give solution for the second equation.
???

Answer by mananth(16946)

3.00 x + 2.00 y = 7.00 .............1
Total value
1,234.00 x + 1,345.00 y = 786.00 .............2
Eliminate y
multiply (1)by 1,345.00
Multiply (2) by -2.00
4,035.00 x 2,690.00 y = 9,415.00
-2,468.00 x -2,690.00 y = -1,572.00
Add the two equations
1,567.00 x = 7,843.00
/ 1,567.00
x = 5.01

plug value of x in (1)
3.00 x + 2.00 y = 7.00
15.02 + 2.00 y = 7.00
2.00 y = 7.00 -15 1/65
2.00 y = -8.02
y = -4.01